How far away can a boy ride his bicycle if he rides away at 10 kilometers per hour and returns at 9 kilometers per hour? The entire trip takes 9.5 hours.
Answer by Alan3354(69443)

How far away can a boy ride his bicycle if he rides away at 10 kilometers per hour and returns at 9 kilometers per hour? The entire trip takes 9.5 hours.
----------------
His avg speed is 2*10*9/(10 + 9) km/hr
= 180/19 km/hr
------
RT distance = r*t = (180/19)*9.5 = 90 km
Each way = 45 km